OpenStudy (anonymous):

What are the characteristics of a radical equation? (1 point) How is solving radical equations similar to solving linear equations? (1 point) Why is it important to check the solutions to a radical equation? (1 point) Create your own radical equation. Describe in complete sentences and demonstrate the process in finding its solution(s). (2 points)

OpenStudy (anonymous):

google it?

OpenStudy (anonymous):

no wait i got it

OpenStudy (anonymous):

Subtract 4 from both sides. Square both the equations (√x+2)^2 = (x - 4)^2 x + 2 = (x - 4)(x - 4) x + 2 = x^2 - 4x - 4x + 16 x + 2 = x^2 - 8x + 16 subtract x ...:Subtract 4 from both sides Square both the equations ( x+2)^2 = (x - 4)^2 x + 2 = (x - 4)(x - 4) x + 2 = x^2 - 4x - 4x + 16 x + 2 = x^2 - 8x + 16 subtract x from both sides 2 = x^2 - 9x + 16 Subtract 2 from both sides 0 = x^2 - 9x + 14 Set up to parentheses with: 0 = (x - 2) (x - 7) set them both equal to 0 x - 2 = 0 x - 7 = 0 x = -2 x = -7 Check your work. The last ones im still doing i have algebra 2 online Question:OK so i have these two equations that i tried solving 10 times and even got help from how they did similar problems in the book but I DID NOT get the same answer as there was in the back of the book! please help! and please just don't don't give me the answer, please show me how to do it with details!! click the link to see the equations: http://s144.photobucket.com/albums/r171/babybella_2007/?action=view&current=eq.jpg In the book of the book , these were the answers the first questions answer is: -5/3 , The second answer is 12/29 I :1. -2 / (x + 3) = 1 / (x + 1) - 2 (x + 1) = 1 (x + 3) -2x - 2 = x + 3 - 2x - x = 3 + 2 - 3x = 5 x = - 5/3 2. (4/x) + (1/3) = 10 4/x = 10 - (1/3) 4/x = (30/3) - (1/3) 4/x = 29/3 4 = (29/3)x 12 = 29x x = 12/29